# Alternating Gradient-Type Algorithm for Bilevel Optimization with Inexact Lower-Level Solutions via Moreau Envelope-based Reformulation

### v1

@@ 1 Introduction
Here, $X\in{\mathbb{R}}^{n}$ and $Y\in{\mathbb{R}}^{m}$ are closed convex sets, $f(x,y):{\mathbb{R}}^{n}\times{\mathbb{R}}^{m}\rightarrow{\mathbb{R}}$ is smooth and convex with respect to $y$ on $Y$ for any $x\in X$, and $g(x,y):{\mathbb{R}}^{n}\times{\mathbb{R}}^{m}\rightarrow{\mathbb{R}}$ is convex but potentially nonsmooth with respect to $y$ on $Y$ for any $x\in X$. Additionally, we assume that $g(x,\cdot):{\mathbb{R}}^{m}\rightarrow{\mathbb{R}}$ is proximal-friendly, enabling efficient evaluation of its proximal operator.
Bilevel optimization originates from economic game theory, particularly the Stackelberg game [1], and has since found diverse applications in economics [2, 3, 4], transportation [5, 6], and machine learning [7, 8, 9, 10, 11]. Notably, many hyperparameter optimization problems can be modeled as bilevel optimization problems with a convex composite lower-level problem as in (2). Bilevel optimization has been extensively studied, with monographs [12, 13, 14, 15] providing a comprehensive overview of its methodologies and applications. For bilevel problems where all defining functions are smooth, the first-order optimality condition is used to reformulate the problem as a mathematical program with equilibrium constraints (MPEC); see, e.g., [16, 17, 18, 19]. However, this approach often encounters difficulties when applied to larger-scale problems or those involving non-smooth lower-level functions, such as those with regularization terms. Smoothing techniques have also been proposed to address nonsmooth issues in specific applications, such as hyperparameter optimization [20, 21]. Another widely used approach for reformulating bilevel optimization problems is based on the value function [22, 23]. Specifically, the bilevel problem is reformulated as the following equivalent problem,
@@
where $v(x):=\inf_{\theta\in Y}\varphi(x,\theta)$ is the value function of the lower-level problem. Under partial calmness condition, Newton-type methods have been developed for this reformulation [24, 25, 26]. For fully convex lower-level problems, (VP) can be treated as a difference-of-convex (DC) program, as studied in [27], which led to applications in bilevel hyperparameter tuning problems [28]. Subsequently, [29] extended these ideas to settings where the fully convex assumption is relaxed, introducing a Moreau envelope-based reformulation and corresponding algorithms.
@@
In recent years, bilevel optimization has garnered significant attention within the machine learning community, driving the development of various methods and algorithms. Among these, gradient-based approaches have emerged as particularly effective due to their simplicity and efficiency [8, 10, 11]. Many of these methods are built upon the computation of hypergradient or its approximation, especially in cases where the lower-level problem is smooth and strongly convex. The strong convexity of the lower-level problem guarantees the uniqueness of its solution, facilitating hypergradient computation [30, 31, 32, 33, 34, 35]. Among these, [34] proposed a two-timescale single-loop stochastic gradient algorithm (TTSA) that efficiently tackles smooth bilevel problems. In addition, recent works have proposed gradient-based approaches based on the value function reformulation (VP) of bilevel problems, see, e.g., [36, 37, 38, 39, 40]. These approaches typically require smoothness conditions for the lower-level problem. Notably, [39] firstly proposed first-order penalty methods for finding $\epsilon$-KKT solutions of (VP) via solving a minimax optimization problem.
@@
A key challenge in gradient-type algorithms derived from the value function reformulation is the computation of the gradient or subgradient of the value function at each iteration [28, 27]. This involves solving the corresponding lower-level problem, even when the gradient exists, as in the case where $\nabla v(x)=\nabla_{x}\varphi(x,y^{*}(x))$, with $y^{*}(x)=\arg\min_{\theta\in Y}\varphi(x,\theta)$ ([41, Remark 4.14]). However, exactly solving the lower-level problem at each iteration can be computationally expensive, potentially limiting the efficiency of the algorithm. In cases where the lower-level problem satisfies uniformly strong convexity or a global PL condition, recent works [37, 38] use inexact solutions to the lower-level problem, replacing the exact solution $y^{*}(x)$ in the gradient computation, constructing an approximation to the value function gradient. But, when the lower-level problem lacks uniform strong convexity and a global PL condition, constructing an approximation to $\nabla v$ using inexact lower-level solutions becomes challenging. Specifically, using an inexact lower-level solution $y_{k}$ with arbitrarily small first-order residual (i.e., $\|\nabla_{y}f(x_{k},y_{k})\|\rightarrow 0$) still results in a fixed gap between the approximate gradient $\nabla_{x}\varphi(x_{k},y_{k})$ and the true gradient $\nabla v(x_{k})$, even as $k\rightarrow\infty$. As an illustrative example, consider the following simple lower-level problem,
@@
where $x,y\in\mathbb{R}$. For $x>0$, $y^{*}(x)=0$, so $v(x)=0$ and $\nabla v(x)=0$. However, for any vanishing sequence $\epsilon_{k}\rightarrow 0$, consider sequences $x_{k}={\epsilon_{k}}/{2}$ and $y_{k}=1$. Here, $y_{k}$ is an inexact solution to (4) for $x=x_{k}$, with a small first-order residual $\|\nabla_{y}\varphi(x_{k},y_{k})\|=\epsilon_{k}$, which can be arbitrarily small as $k\rightarrow\infty$. Despite this, the gradient approximation $\nabla_{x}\varphi(x_{k},y_{k})$ will still have a constant gap to the true gradient $\nabla v(x_{k})$, since $\|\nabla_{x}\varphi(x_{k},y_{k})-\nabla v(x_{k})\|=1$ even as $k\rightarrow\infty$. This example demonstrates the difficulty of constructing accurate value function gradient approximations and developing gradient-type algorithms with inexact solutions to the lower-level problem for bilevel optimization when the lower-level problem does not have uniformly strong convexity.
@@
where $\gamma>0$. When the lower-level problem is convex with respect to $y$, this reformulation is equivalent to the original bilevel optimization problem (1) (see [29, Theorem 1]). Directional optimality conditions for $({\rm VP})_{\gamma}$ are discussed in [42]. However, as shown in [29] and [27, Proposition 7], classical constraint qualification (CQ) conditions do not hold at any feasible point of $({\rm VP})_{\gamma}$, motivating us to study the following relaxed approximation problem,
@@
where $\epsilon>0$. [29, Proposition 6] shows that for any $\epsilon_{0}>0$, there exists $\epsilon>0$ such that a local minimizer of $({\rm VP})_{\gamma}^{\epsilon}$ is $\epsilon_{0}$-close to the solution set of the original bilevel optimization problem (1). In [29], a double-loop difference-of-convex algorithm was developed to solve $({\rm VP})_{\gamma}^{\epsilon}$ for general bilevel optimization problems. This algorithm requires solving two optimization subproblems at each iteration, including the proximal lower-level problem (6). Several subsequent works [43, 44, 45] have adopted ideas akin to the Moreau envelope-based reformulation to develop single-loop gradient-based bilevel algorithms. However, these works limit their convergence analysis to the decreasing property of certain merit functions, lacking rigorous theoretical support for the convergence of iterates.
@@
  • We propose AGILS, which is an alternating gradient-type algorithm with an adaptive penalty parameter update strategy for solving the approximation bilevel optimization problem $({\rm VP})_{\gamma}^{\epsilon}$. A key feature of AGILS is the introduction of a verifiable inexact criterion for the proximal lower-level problem (6). Unlike the double-loop difference-of-convex algorithm developed in [29], which requires exact solution to the proximal lower-level problem at each iteration, AGILS allows for inexact solutions, enhancing both the flexibility and computational efficiency of the algorithm. In particular, the inexact criterion also facilitates the integration of various efficient methods for solving convex composite models within the proximal lower-level problem.
@@
  • The proposed AGILS can efficiently handle high-dimensional bilevel optimization problems, benefiting from its use of only an inexact solution to the proximal lower-level problem (6) and an alternating gradient-type update scheme. To evaluate its performance, we conduct numerical experiments on a toy example and the sparse group Lasso bilevel hyperparameter selection problem. The numerical results demonstrate the efficiency of the AGILS compared to other commonly used approaches.
@@ 2.2 Basic Assumptions
Smoothness and weak convexity conditions are required for the lower-level objective functions. We recall that a function $h:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}$ is $(\rho_{1},\rho_{2})$-weakly convex on a convex set $X\times Y$ if $h(x,y)+\frac{\rho_{1}}{2}\|x\|^{2}+\frac{\rho_{2}}{2}\|y\|^{2}$ is convex on $X\times Y$. This notion generalizes convexity ($\rho_{1}=\rho_{2}=0$). Weakly convex functions [46], extend the class of convex functions to include many nonconvex functions. For weakly convex functions, the Fréchet (regular) subdifferential coincides with both the limiting subdifferential and the Clarke subdifferential (see, e.g., [47, Proposition 3.1 and Theorem 3.6]).
@@
Assumption 3(1) implies $(\rho_{f_{1}},\rho_{f_{2}})$-weak convexity with $\rho_{f_{1}}=\rho_{f_{2}}=L_{f}$ of $f(x,y)$ [48, Lemma 5.7]. Furthermore, the lower-level objective $\varphi(x,y)$ is $(\rho_{\varphi_{1}},\rho_{\varphi_{2}})$-weakly convex on $X\times Y$, where $\rho_{\varphi_{1}}=\rho_{f_{1}}+\rho_{g_{1}}$ and $\rho_{\varphi_{2}}=\rho_{f_{2}}+\rho_{g_{2}}$.
According to [49, Lemma 4.2], $g(x,y)=h(c(x,y))$ is $L_{h}L_{c}$-weakly convex if $h:\mathbb{R}^{d}\to\mathbb{R}$ is a closed convex $L_{h}$-Lipschitz continuous function and $c:\mathbb{R}^{n+m}\to\mathbb{R}^{d}$ is a smooth function with an $L_{c}$-Lipschitz continuous Jacobian. Below, we provide more specific examples of $g(x,y)$ satisfying 3.
Lemma 1.
Let $g(x,y)$ satisfy one of the following conditions. Then $g(x,y)$ satisfies Assumption 3.
  • (1)
    $g(x,y)=x\|y\|_{1}$. Then $g(x,y)$ is $(m,1)$-weakly convex on +×$\mathbb{R}_{+}\times\mathbb{R}^{m}$.
  • (2)
    $g(x,y)=\sum_{j=1}^{J}x_{j}\|y^{(j)}\|_{2}$, where $y^{(j)}$ denotes the $j$-th group of $y$. Then $g(x,y)$ is $(1,1)$-weakly convex on +$\mathbb{R}^{J}_{+}\times\mathbb{R}^{m}$.
  • (3)
    $g(x,y)=xP(y)$, where $P(y)$ is convex and $L_{P}$-Lipschitz continuous on $Y$. Then $g(x,y)$ is $(\rho_{g_{1}},\rho_{g_{2}})$-weakly convex on +×$\mathbb{R}_{+}\times\mathbb{R}^{m}$ for any $\rho_{g_{1}},\rho_{g_{2}}\geq 0$ satisfying $\rho_{g_{1}}\rho_{g_{2}}\geq L_{P}^{2}$.
Proof.
For (1), the proof follows a similar discussion as in Section 6.1 of [29], using a different estimation of weak convexity constants. We note that $x\|y\|_{1}+\frac{m}{2}x^{2}+\frac{1}{2}\|y\|^{2}_{2}=\sum_{i=1}^{m}\frac{1}{2}\left(x+|y_{i}|\right)^{2}$, which is the composition of a linear function with a convex function. By [50, Proposition 1.54], the sum is convex, implying $x\|y\|_{1}$ is $(m,1)$-weakly convex. For (2), see Section 6.2 of [29]. To establish (3), we recall that $g(x,y)$ is $(\rho_{g_{1}},\rho_{g_{2}})$-weakly convex on $X\times Y$ if and only if, for any points $(x_{1},y_{1}),(x_{2},y_{2})\in X\times Y$ and $a\in[0,1]$, the approximate secant inequality holds: $g(ax_{1}+(1-a)x_{2},ay_{1}+(1-a)y_{2})\leq ag(x_{1},y_{1})+(1-a)g(x_{2},y_{2})+a(1-a)/2\cdot(\rho_{g_{1}}\|x_{1}-x_{2}\|^{2}+\rho_{g_{2}}\|y_{1}-y_{2}\|^{2})$. Now, consider $g(x,y)=xP(y)$. For any points $(x_{1},y_{1}),(x_{2},y_{2})\in\mathbb{R}_{+}\times\mathbb{R}^{m}$ and $a\in[0,1]$, we have
$g(ax_{1}+(1-a)x_{2},ay_{1}+(1-a)y_{2})-ag(x_{1},y_{1})+(1-a)g(x_{2},y_{2})$
@@
$\leq$ $a(1-a)L_{P}\|x_{1}-x_{2}\|\left\|y_{1}-y_{2}\right\|.$
where the first inequality follows from the convexity of $P(y)$. Therefore, when $\rho_{g_{1}}\rho_{g_{2}}\geq L_{P}^{2}$, the approximate secant inequality is satisfied. This establishes (3). ∎

### v2

@@ 1 Introduction
Here, $X\subset{\mathbb{R}}^{n}$ and $Y\subset{\mathbb{R}}^{m}$ are closed convex sets, $f(x,y):{\mathbb{R}}^{n}\times{\mathbb{R}}^{m}\rightarrow{\mathbb{R}}$ is smooth and convex with respect to $y$ on $Y$ for any $x\in X$, and $g(x,y):{\mathbb{R}}^{n}\times{\mathbb{R}}^{m}\rightarrow{\mathbb{R}}$ is convex but potentially nonsmooth with respect to $y$ on $Y$ for any $x\in X$. Additionally, we assume that $g(x,\cdot):{\mathbb{R}}^{m}\rightarrow{\mathbb{R}}$ is proximal-friendly, enabling efficient evaluation of its proximal operator.
Bilevel optimization originates from economic game theory, particularly the Stackelberg game [57], and has since found diverse applications in economics [16, 32, 47], transportation [19, 46], and machine learning [8, 29, 38, 41, 63]. Notably, many hyperparameter optimization problems can be modeled as bilevel optimization problems with a convex composite lower-level problem as in (2). Bilevel optimization has been extensively studied, with monographs [18, 20, 22, 21] providing a comprehensive overview of its methodologies and applications. For bilevel problems where all defining functions are smooth, the first-order optimality conditions are used to reformulate the problem as a mathematical program with equilibrium constraints (MPEC); see, e.g., [2, 6, 45, 54]. However, this approach often encounters difficulties when applied to larger-scale problems or those involving non-smooth lower-level functions, such as those with regularization terms. Smoothing techniques have also been proposed to address nonsmooth issues in specific applications, such as hyperparameter optimization [1, 52]. Another widely used approach for reformulating bilevel optimization problems is based on the value function [53, 62]. Specifically, the bilevel problem is reformulated as the following equivalent problem,
@@
where $v(x):=\inf_{\theta\in Y}\varphi(x,\theta)$ is the value function of the lower-level problem. Under the partial calmness condition, Newton-type methods have been developed for this reformulation [27, 28, 36]. For fully convex lower-level problems, (VP) can be treated as a difference-of-convex (DC) program, as studied in [61], which led to applications in bilevel hyperparameter tuning problems [30]. Subsequently, [31] extended these ideas to settings where the fully convex assumption is relaxed, introducing a Moreau envelope-based reformulation and corresponding algorithms.
@@
In recent years, bilevel optimization has garnered significant attention within the machine learning community, driving the development of various methods and algorithms. Among these, gradient-based approaches have emerged as particularly effective due to their simplicity and efficiency [29, 41, 63]. Many of these methods are built upon the computation of hypergradient or its approximation, especially in cases where the lower-level problem is smooth and strongly convex. The strong convexity of the lower-level problem guarantees the uniqueness of its solution, facilitating hypergradient computation [10, 17, 26, 33, 34, 35]. Among these, [34] proposed a two-timescale single-loop stochastic gradient algorithm (TTSA) that efficiently tackles smooth bilevel problems. In addition, recent works have proposed gradient-based approaches based on the value function reformulation (VP) of bilevel problems, see, e.g., [39, 40, 43, 44, 55]. These approaches typically require smoothness conditions for the lower-level problem. Notably, [44] first proposed first-order penalty methods for finding $\epsilon$-KKT solutions of (VP) via solving a minimax optimization problem.
@@
A key challenge in gradient-type algorithms derived from the value function reformulation is the computation of the gradient or subgradient of the value function at each iteration [30, 61]. This involves solving the corresponding lower-level problem, even when the gradient exists, as in the case where $\nabla v(x)=\nabla_{x}\varphi(x,y^{*}(x))$, with $y^{*}(x)=\arg\min_{\theta\in Y}\varphi(x,\theta)$ ([15, Remark 4.14]). However, exactly solving the lower-level problem at each iteration can be computationally expensive, potentially limiting the efficiency of the algorithm. In cases where the lower-level problem satisfies uniformly strong convexity or a global PL condition, recent works [40, 43] use inexact solutions to the lower-level problem, replacing the exact solution $y^{*}(x)$ in the gradient computation, constructing an approximation to the value function gradient. But, when the lower-level problem lacks uniform strong convexity or a global PL condition, constructing an approximation to $\nabla v$ using inexact lower-level solutions becomes challenging. Specifically, using an inexact lower-level solution $y_{k}$ with arbitrarily small first-order residual (i.e., $\|\nabla_{y}\varphi(x_{k},y_{k})\|\rightarrow 0$) can still result in a fixed gap between the approximate gradient $\nabla_{x}\varphi(x_{k},y_{k})$ and the true gradient $\nabla v(x_{k})$, even as $k\rightarrow\infty$. As an illustrative example, consider the following simple lower-level problem,
@@
where $x,y\in\mathbb{R}$. For $x>0$, $y^{*}(x)=0$, so $v(x)=0$ and $\nabla v(x)=0$. However, for any vanishing sequence $\epsilon_{k}\rightarrow 0$, consider sequences $x_{k}={\epsilon_{k}}/{2}$ and $y_{k}=1$. Here, $y_{k}$ is an inexact stationary point to (4) for $x=x_{k}$, with a small first-order residual $\|\nabla_{y}\varphi(x_{k},y_{k})\|=\epsilon_{k}$, which can be arbitrarily small as $k\rightarrow\infty$. Despite this, the gradient approximation $\nabla_{x}\varphi(x_{k},y_{k})$ will still have a constant gap to the true gradient $\nabla v(x_{k})$, since $\|\nabla_{x}\varphi(x_{k},y_{k})-\nabla v(x_{k})\|=1$ even as $k\rightarrow\infty$. This example demonstrates the difficulty of constructing accurate value function gradient approximations and developing gradient-type algorithms with inexact solutions to the lower-level problem for bilevel optimization when the lower-level problem does not have uniformly strong convexity.
@@
where $\gamma>0$. When the lower-level problem is convex with respect to $y$, this reformulation is equivalent to the original bilevel optimization problem (1) (see [31, Theorem 1]). Directional optimality conditions for $({\rm VP})_{\gamma}$ are discussed in [5]. However, as shown in [31] and [61, Proposition 7], classical constraint qualification (CQ) conditions do not hold at any feasible point of $({\rm VP})_{\gamma}$, motivating us to study the following relaxed approximation problem,
@@
where $\epsilon>0$. [31, Proposition 6] shows that for any $\epsilon_{0}>0$, there exists $\epsilon>0$ such that a local minimizer of $({\rm VP})_{\gamma}^{\epsilon}$ is $\epsilon_{0}$-close to the solution set of the original bilevel optimization problem (1). In [31], a double-loop difference-of-convex algorithm was developed to solve $({\rm VP})_{\gamma}^{\epsilon}$ for general bilevel optimization problems. This algorithm requires solving two optimization subproblems at each iteration, including the proximal lower-level problem (6). Several subsequent works [42, 59, 60] have adopted ideas akin to the Moreau envelope-based reformulation to develop single-loop gradient-based bilevel algorithms. However, these works limit their convergence analysis to the decreasing property of certain merit functions, lacking rigorous theoretical support for the convergence of iterates.
@@
  • We propose AGILS, an alternating gradient-type algorithm designed to solve the approximate bilevel optimization problem $({\rm VP})_{\gamma}^{\epsilon}$. A key feature of AGILS is the incorporation of a verifiable inexactness criterion for the proximal lower-level problem (6). Unlike the double-loop difference-of-convex algorithm developed in [31], which requires an exact solution to the proximal lower-level problem at each iteration, AGILS allows for inexact solutions, enhancing both the flexibility and computational efficiency of the algorithm. Furthermore, to ensure the feasibility of the iterates, we incorporate an adaptive penalty parameter update strategy and a feasibility correction procedure into AGILS.
@@
  • The proposed AGILS can efficiently handle high-dimensional bilevel optimization problems, benefiting from its use of only an inexact solution to the proximal lower-level problem (6) and an alternating gradient-type update strategy. This strategy decouples the nonsmoothness of $g(x,y)$, enabling AGILS to handle nonsmooth terms effectively. To evaluate its performance, we conduct numerical experiments on a toy example and the sparse group Lasso bilevel hyperparameter selection problem. The numerical results demonstrate the efficiency of the AGILS compared to other commonly used approaches.
@@ 2.2 Basic Assumptions
Smoothness and weak convexity conditions are required for the lower-level objective functions. We recall that a function $h:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}$ is $(\rho_{1},\rho_{2})$-weakly convex on a convex set $X\times Y$ if $h(x,y)+\frac{\rho_{1}}{2}\|x\|^{2}+\frac{\rho_{2}}{2}\|y\|^{2}$ is convex on $X\times Y$. This notion generalizes convexity ($\rho_{1}=\rho_{2}=0$). Weakly convex functions [51], extends the class of convex functions to include many nonconvex functions. For weakly convex functions, the Fréchet (regular) subdifferential coincides with both the limiting subdifferential and the Clarke subdifferential (see, e.g., [50, Proposition 3.1 and Theorem 3.6]).
@@
Assumption 3(1) implies $(\rho_{f_{1}},\rho_{f_{2}})$-weak convexity with $\rho_{f_{1}}=\rho_{f_{2}}=L_{f}$ of $f(x,y)$ [7, Lemma 5.7]. Furthermore, the lower-level objective $\varphi(x,y)$ is $(\rho_{\varphi_{1}},\rho_{\varphi_{2}})$-weakly convex on $X\times Y$, where $\rho_{\varphi_{1}}=\rho_{f_{1}}+\rho_{g_{1}}$ and $\rho_{\varphi_{2}}=\rho_{f_{2}}+\rho_{g_{2}}$.
According to [24, Lemma 4.2], $g(x,y)=h(c(x,y))$ is $L_{h}L_{c}$-weakly convex if $h:\mathbb{R}^{d}\to\mathbb{R}$ is a closed convex $L_{h}$-Lipschitz continuous function and $c:\mathbb{R}^{n+m}\to\mathbb{R}^{d}$ is a smooth function with an $L_{c}$-Lipschitz continuous Jacobian. Below, we provide more specific examples of $g(x,y)$ satisfying Assumption 3.
Example 1.
Consider $g(x,y)=xP(y)$, where $P(y)$ is a convex and $L_{P}$-Lipschitz continuous function on $Y$. Then $g(x,y)$ is $(\rho_{g_{1}},\rho_{g_{2}})$-weakly convex on +×$\mathbb{R}_{+}\times\mathbb{R}^{m}$ for any $\rho_{g_{1}},\rho_{g_{2}}\geq 0$ satisfying $\rho_{g_{1}}\rho_{g_{2}}\geq L_{P}^{2}$.
To verify this, recall that $g(x,y)$ is $(\rho_{g_{1}},\rho_{g_{2}})$-weakly convex on $X\times Y$ if and only if, for any points $(x_{1},y_{1}),(x_{2},y_{2})\in X\times Y$ and $a\in[0,1]$, the approximate secant inequality holds: $g(ax_{1}+(1-a)x_{2},ay_{1}+(1-a)y_{2})\leq ag(x_{1},y_{1})+(1-a)g(x_{2},y_{2})+a(1-a)/2\cdot(\rho_{g_{1}}\|x_{1}-x_{2}\|^{2}+\rho_{g_{2}}\|y_{1}-y_{2}\|^{2})$. Now, consider $g(x,y)=xP(y)$. For any points $(x_{1},y_{1}),(x_{2},y_{2})\in\mathbb{R}_{+}\times\mathbb{R}^{m}$ and $a\in[0,1]$, we have
$g(ax_{1}+(1-a)x_{2},ay_{1}+(1-a)y_{2})-ag(x_{1},y_{1})-(1-a)g(x_{2},y_{2})$
@@
$\leq$ $a(1-a)L_{P}\|x_{1}-x_{2}\|\left\|y_{1}-y_{2}\right\|,$
where the first inequality follows from the convexity of $P(y)$. Therefore, when $\rho_{g_{1}}\rho_{g_{2}}\geq L_{P}^{2}$, the approximate secant inequality is satisfied, proving the claim.
Next, we present two practical instances of this form, LASSO and group LASSO:
Example 2.
The following structured regularizers satisfy the above conditions:
(1) Let $g(x,y)=x\|y\|_{1}$. Then $g(x,y)=xP(y)$ with $P(y)=\|y\|_{1}$, which is $\sqrt{m}$-Lipschitz on $\mathbb{R}^{m}$. Thus, $g(x,y)$ is $(m,1)$-weakly convex on +×$\mathbb{R}_{+}\times\mathbb{R}^{m}$.
(2) Let $g(x,y)=\sum_{j=1}^{J}x_{j}\|y^{(j)}\|_{2}$, where $y^{(j)}$ denotes the $j$-th group of $y$. Then $g(x,y)=\sum_{j=1}^{J}x_{j}P_{j}(y^{(j)})$ with $P_{j}(y^{(j)})=\|y^{(j)}\|_{2}$, which is $1$-Lipschitz. Thus, $g(x,y)$ is $(1,1)$-weakly convex on +$\mathbb{R}^{J}_{+}\times\mathbb{R}^{m}$.